OpenStudy (anonymous):

Find the shortest distance from the point (-8,-6) to the line y=5/3x - 4.

OpenStudy (anonymous):

Find the line perpendicular to y=(5/3)x - 4 that passes through (-8,-6). Hint the slope of this perpendicular line will be the negative inverse of the slope of your original line (i.e. -1/m where m is the slope of y=(5/3)x - 4). See if this helps